# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v4

@@ 7.1 Effect of the Expected Photon Numbers
7.1.1 Effect of the Exected Photon Number on Cost
Since $\mu$ is the expected number of photons in unit time, this can be reduced by using heavy attenuation. This will directly affect the cost. Heavy attenuation requires a large number of high-quality attenuators, increasing the production cost for the random numbers. Thus
@@ 7.1.1 Effect of the Exected Photon Number on Cost
7.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (6.2), the probability, $P(0)$, of no photon detection within time $T$ is given by
@@ 7.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
Figure 8: Plot shows how random number generation rate (bit per unit time) changes with the expected photon number per unit time ($\mu$). Five different plots have been shown for five different values of detection efficiency ($d$). From the plot, it can be easily seen that if $\mu$ decreases, the random number generation rate of the QRNG will also decrease.
Suppose, for some $\mu$, the expected number of external cycles to produce a random number of length $k$, where each cycle contains $N$ time-bins, is given by $N_{c}$. Then $N_{c}=\langle X\rangle$, where $X$ is a random variable, whose probability distribution is negative binomial distribution, given by
@@
for $x=k^{\prime},k^{\prime}+1,k^{\prime}+2,\dots$, where $k^{\prime}=\lceil\frac{k}{\log_{2}N}\rceil$ and $P_{\mu}(0)=e^{-\mu Td}$. Therefore,
$N_{c}=\langle X\rangle=\frac{k^{\prime}P(0)}{1-P(0)}=\frac{k^{\prime}}{e^{\mu Td}-1}.$ (45)
@@
$r=\frac{k^{\prime}}{TN_{c}}=\frac{e^{\mu Td}-1}{T}\text{ bits per unit time}.$ (46)
@@
$r\approx\mu d\text{ bits per unit time}.$ (47)
@@
7.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (32) of the Subsection 6.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 9. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $[(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $[(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as $(i-1)$-th bin. Similarly, if the actual detection happens within time interval $[i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
@@ 7.2 Effect of the timing error of the device on the External Reference Cycle
7.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
@@
7.4 Effect of Detection Efficiency
@@
7.4.1 Effect of the Detection Efficiency on Cost
@@
small value of $\text{small value of }d\implies\begin{cases}\text{high randomness}\\ \text{high cost}\end{cases}$ (49)
7.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
@@
Figure 10: Plot shows how random number generation rate (bit per unit time) changes with the detection efficiency ($d$). Five different plots have been shown for five different values of $\mu$, the expected photon number per unit time. From the plot, it can be easily seen that if $d$ decreases, the random number generation rate of the QRNG will also decrease.
@@
Although a bad detector with a small detection efficiency, $d$, reduces production costs, this leads to a high chance of no photon detection in a particular external reference cycle. From Eq. (47) it can be easily seen that if $T$ decreases, the rate of generating random numbers will also decrease. The graph in Fig. 10 shows how this rate changes with $\mu$ for fixed $T$ and $d$. Therefore, it is not possible to choose a detector with arbitrarily small detection efficiency.
7.5 Performance and Cost of the QRNG
@@
Figure 11: Plot showing relation between random number generation rate, $r$, expected photon number in unit time, $\mu$, and detection efficiency, $d$. If the values of $\mu$ and $d$ are small, the generation rate is also small. Therefore, to generate random numbers with high rate, we have to choose both $\mu$ and $d$ large.
Figure 12: Heat map denoting how (a) randomness, (b) random number generation rate and (c) cost varies depending on the expected photon number per unit time, $\mu$, external reference cycle, $T$ and the detection efficiency, $d$. Figure (a) shows that we must keep the product $\mu Td$ small to get maximum randomness. However, (b) shows if the $\mu d$ is small, the generation rate will be minimal. On the other hand, from (c), low-cost demands a small value of $d$ and a high value of $\mu$.
@@
Figure 11 is a surface plot of how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (32) of Subsection 6.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Therefore, this needs some optimization while choosing $\mu$ and $d$.
@@
The above discussion has been described as a heat map in Fig. 12. In the figure, dark shade denotes low value and light shade denotes high value of randomness, generation rate and cost in Figures 12(a), 12(b) and 12(c) respectively. Figure 12(a) shows that high randomness demands a small value of the product $\mu Td$. However, the product $\mu d$ cannot be very small as Fig. 12(b) indicates. A small $\mu d$ would result in a low random number generation rate. On the other hand, to keep the cost of the QRNG low, we need a small $d$ but a large $\mu$.
@@
8 Results of Randomness Tests
Figure 13: NIST result. Bar charts are plotted for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$, (i) $\mu Td=0.45$, (j) $\mu Td=0.50$. For each $\mu Td$, three $1$ GB files are generated by simulating probability distribution (31), and the NIST tests have been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value $=0.01$
Random numbers and their randomness testing have been discussed in Appendix E. The randomness of the random numbers is tested using statistical tests. Here we used four common statistical testing suites namely NIST, Dieharder, AIS-31 and ENT.
@@ 8 Results of Randomness Tests
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section 5, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
@@ 9 Discussion
Acknowledgement
S. Das would like to acknowledge the support received by the Quantum Delta NL KAT-2 project.
@@ Appendix A A Comparison between trusted QRNG and DI-QRNG
A device-independent quantum protocol uses quantum entanglement to ensure security. This can be verified using Bell inequality, quantum discord, etc. [71, 84, 85, 86]. On the other hand, if a protocol relies on single photons as a resource, the absence of entanglement prevents the application of the validation procedures mentioned above. However, we can still verify its quantumness using photon-count statistics [52, 53, 89, 90], discussed in Section  2 and Appendix B, to improve the security. A detailed comparison indicating advantages and disadvantages is given in Table 2.
@@
2. Security relies on trusted assumptions. Hence, less secure. 2. Security relies on Bell violation or similar quantum principles [71, 84, 85, 86]. Hence, more secure.
@@
2. No Bell test required. 2. Loophole-free Bell test is needed and performing a loophole-free Bell test is difficult [102, 103, 104, 105, 106].
3. Can be produced in the very compact form of USB or chip. 3. To avoid locality loophole in the Bell test, two entangled particles must be space-like separated [107, 108]. This forces the device to be either large or trusted during the Bell test.
4. Compatible for portable devices. 4. Incompatible in use cases with portable devices, e.g., automobiles, satellites, point of sale (POS) terminals, portable phones, laptops, etc. [36]
@@
The second-order correlation function [53] of light is defined by
@@
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 00, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes [53]:
@@
The relation between photon count distribution and $g^{2}(\tau)$ is given by [52]
@@
In the above equation, if $g^{2}(\tau)\leq 1$ for all $\tau$, it implies $\langle(\Delta n)^{2}\rangle<\langle N\rangle$, i.e., a sub-Poissonian statistics, indicating non-classicality of light. However, if $g^{2}(\tau)\leq 1$ for some $\tau$, there is no direct relation between the sub-Poissonian photon count statistic and the antibunching of photons. Xou and Mandel [52] showed that if a source emits photons in two different nonvacuum modes, both having occupation number $n/2$, but different frequencies $\omega_{1}$ and $\omega_{2}$, the second-order correlation function takes the form

### v5

@@ VII.1 Effect of the Expected Photon Numbers
VII.1.1 Effect of the Exected Photon Number on Cost
Since $\mu$ is the expected number of photons in unit time, this can be reduced by using heavy attenuation. This will directly affect the cost. Heavy attenuation requires a large number of attenuators, increasing the production cost for the random numbers. Thus
@@ VII.1.1 Effect of the Exected Photon Number on Cost
VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (VI.2), the probability, $P(0)$, of no photon detection within time $T$ is given by
@@ VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
Suppose, for some $\mu$, the expected number of external cycles to produce a random number of length $k$ bits, where each cycle contains $N$ time-bins, is given by $N_{c}$. Then $N_{c}=\langle X\rangle$, where $X$ is a random variable, whose probability distribution is negative binomial distribution, given by
@@
for $x=k^{\prime},k^{\prime}+1,k^{\prime}+2,\dots$, where $k^{\prime}=\lceil\frac{k}{\log N}\rceil$ is the minimum number of cycles required to produce $k$-bit random number and $P_{\mu}(0)=e^{-\mu Td}$. Therefore,
$N_{c}=\langle X\rangle=\frac{k^{\prime}}{1-P(0)}=\frac{k^{\prime}}{1-e^{-\mu Td}}.$ (45)
@@
$r=\frac{k}{TN_{c}}\approx\frac{1-e^{-\mu Td}}{T}\log N\text{ bits per unit time}.$ (46)
@@
$r\approx\mu d\log N\text{ bits per unit time}.$ (47)
Figure 8: Plot shows how random number generation rate (bit per unit time) changes with the expected photon number per unit time ($\mu$) given by Eq. (47) when $d$ is constant. Five different plots have been shown for five different values of detection efficiency ($d$). From the plot, it can be easily seen that if $\mu$ decreases, the random number generation rate of the QRNG will also decrease.
@@
VII.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (32) of the Subsection VI.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 9. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $[(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $[(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as $(i-1)$-th bin. Similarly, if the actual detection happens within time interval $[i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
@@ VII.2 Effect of the timing error of the device on the External Reference Cycle
VII.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
@@
VII.4 Effect of Detection Efficiency
@@
VII.4.1 Effect of the Detection Efficiency on Cost
@@
small value of $\text{small value of }d\implies\begin{cases}\text{high randomness}\\ \text{low cost}\end{cases}$ (49)
VII.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
@@
Figure 10: Plot shows how random number generation rate (bit per unit time) changes with the detection efficiency ($d$) given by Eq. (47) when $\mu$ is constant. Five different plots have been shown for five different values of $\mu$, the expected photon number per unit time. From the plot, it can be easily seen that if $d$ decreases, the random number generation rate of the QRNG will also decrease.
@@
Although a bad detector with a small detection efficiency, $d$, reduces production costs, this leads to a high chance of no photon detection in a particular external reference cycle. From Eq. (47) it can be easily seen that if $T$ decreases, the rate of generating random numbers will also decrease. The graph in Fig. 10 shows how this rate changes with $d$ for fixed $T$ and $\mu$. Therefore, it is not possible to choose a detector with arbitrarily small detection efficiency.
VIII Performance and Cost of the QRNG
@@
Figure 11: Plot showing relation between random number generation rate, $r$, expected photon number in unit time, $\mu$, and detection efficiency, $d$ given by Eq. (47). If the values of $\mu$ and $d$ are small, the generation rate is also small. Therefore, to generate random numbers with high rate, we have to choose both $\mu$ and $d$ large.
@@
We can see from Eq. (47) that random number generation rate $r\approx\mu d\log N$, where $\mu$ is the expected number of photons in unit time, $d$ is the efficiency of detector and $N$ is the number of time-bins. Figure 11 is a surface plot of Eq. (36) to show how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (32) of Subsection VI.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Also, from Eq. (42)and (49), we can write the cost of the QRNG as
$\text{cost}=\frac{\alpha}{\mu}+\beta d,$ (50)
for some positive constants $\alpha$ and $\beta$. Therefore, this needs some optimization while choosing $\mu$ and $d$.
Considering all the discussions in the previous section, we can now write random number generation rate $r$ as a function of randomness. From Eq. (31), (36) and (47), we have
$2\epsilon$ $=\sum_{i=1}^{N}\left|\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}-\frac{1}{N}\right|$
$=\sum_{i\leq N/2}\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}-\frac{1}{N}\right)+\sum_{i>N/2}\left(\frac{1}{N}-\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}\right)$
$=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left(\sum_{i\leq N/2}e^{-i\frac{\mu Td}{N}}-\sum_{i>N/2}e^{-i\frac{\mu Td}{N}}\right)$
$=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left(\frac{e^{-\frac{\mu Td}{N}}\left(e^{-\frac{\mu Td}{2}}-1\right)}{e^{-\frac{\mu Td}{N}}-1}-\frac{e^{-\frac{\mu Td}{N}\left(\frac{N}{2}+1\right)}\left(e^{-\frac{\mu Td}{2}}-1\right)}{e^{-\frac{\mu Td}{N}}-1}\right)$
$=\frac{\left(1-e^{-\frac{\mu Td}{2}}\right)^{2}}{1-e^{-\mu Td}}=\frac{1-e^{-\frac{\mu Td}{2}}}{1+e^{-\frac{\mu Td}{2}}}.$ (51)
Therefore we can write,
$2\epsilon\left(1+e^{-\frac{\mu Td}{2}}\right)=1-e^{-\frac{\mu Td}{2}}$
$\implies$ $e^{\frac{\mu Td}{2}}=\frac{1+2\epsilon}{1-2\epsilon}$
$\implies$ $r=\frac{2\log N}{T}\ln{\frac{1+2\epsilon}{1-2\epsilon}}.$ (52)
Figure 12: Maximum rate is plotted against $\epsilon$ as Eq. (53) with $T_{min}=Nk_{tol}\delta_{t}=1$ unit of time. It shows that the maximum rate increases very quickly with $\epsilon$. Since from Definition 1, randomness decreases with the increment of $\epsilon$, the maximum rate decreases as randomness increases.
Equation (52) gives the random number generation rate when the generated number is $\epsilon$-random, according to Definition 1. Therefore using Eq. (48), we can write the maximum rate for some given $\epsilon$ as
[…]
where $N$ is the number of time-bins, $\delta_{t}$ is the maximum timing error in photon detection and $k_{tol}$ is a parameter corresponding to the tolerance of this timing error. Fig. 12 plots this maximum rate against $\epsilon$. It shows that the rate decreases with the increment of randomness.
Also, we can write the minimum entropy (41) as a function of $\epsilon$ as
$H_{min}=\log\frac{8\epsilon}{(1+2\epsilon)^{2}\left(1-\left(\frac{1-2\epsilon}{1+2\epsilon}\right)^{\frac{2}{N}}\right)}$ (54)
In fig. 13 we plot this minimum entropy against $\epsilon$.
Figure 13: Minimum entropy, Eq. (54), is plotted against $\epsilon$ for $N=256$. It shows that the minimum entropy decreases, which is a clear indication of decrement of randomness, as $\epsilon$ increases.
@@
Figure 14: Heat map denoting how (a) $\epsilon$-randomness, (b) minimum entropy, (c) random number generation rate and (d) cost varies depending on the expected photon number per unit time, $\mu$, external reference cycle, $T$ and the detection efficiency, $d$. Figure (a) and (b) show the $\epsilon$-randomness in Eq. (51) and minimum entropy in Eq. (41), respectively, indicating that we must keep the product $\mu Td$ small to get high randomness. However, (b) for Eq. (47), with $N=256$, shows if the $\mu d$ is small, the generation rate will be minimal. On the other hand, from (c) for Eq. (50) with $\alpha=1/20,\beta=40$, low-cost demands a small value of $d$ and a high value of $\mu$.
The above discussion has been described as a heat map in Fig. 14. In the figure, dark shade denotes low value and light shade denotes high value of $\epsilon$, minimum entropy, generation rate and cost in Figures 14(a), 14(b), 14(c) and 14(d), respectively. Figure 14(a) and (b) shows that high randomness demands a small value of the product $\mu Td$. However, the product $\mu d$ cannot be very small as Fig. 14(b) indicates. A small $\mu d$ would result in a low random number generation rate. On the other hand, to keep the cost of the QRNG low, we need a small $d$ but a large $\mu$.
@@
IX Results of Randomness Tests
Figure 15: NIST result. $p$-values for passed tests are plotted as bar charts for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$ , (i) $\mu Td=0.45$, (j) $\mu Td=0.50$. For each $\mu Td$, three $1$ GB files are generated by simulating probability distribution (31), and the NIST tests have been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value $=0.01$
Random numbers and their randomness testing have been discussed in E. The randomness of the random numbers is tested using statistical tests. Here we used four common statistical testing suites namely NIST, Dieharder, AIS-31 and ENT.
@@ IX Results of Randomness Tests
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section V, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
@@ Appendix A A Comparison between trusted QRNG and DI-QRNG
A device-independent quantum protocol uses quantum entanglement to ensure security. This can be verified using Bell inequality, quantum discord, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . On the other hand, if a protocol relies on single photons as a resource, the absence of entanglement prevents the application of the validation procedures mentioned above. However, we can still verify its quantumness using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 , discussed in Section  II and B, to improve the security. A detailed comparison indicating advantages and disadvantages is given in Table 2.
@@
2. Security relies on trusted assumptions. Hence, less secure. 2. Security relies on Bell violation or similar quantum principles acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . Hence, more secure.
@@
2. No Bell test required. 2. Loophole-free Bell test is needed and performing a loophole-free Bell test is difficult PhysRevA.47.R747 ; GISIN1999103 ; PhysRevA.66.042111 ; PhysRevA.81.032106 ; Larsson_2014 .
3. Can be produced in the very compact form of USB or chip. 3. To avoid locality loophole in the Bell test, two entangled particles must be space-like separated PhysRevLett.115.250402 ; storz2023loophole . This forces the device to be either large or trusted during the Bell test.
4. Compatible for portable devices. 4. Incompatible in use cases with portable devices, e.g., automobiles, satellites, point of sale (POS) terminals, portable phones, laptops, etc. quantis
@@
The second-order correlation function fox2006quantum of light is defined by
@@
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 00, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes fox2006quantum :
@@
The relation between photon count distribution and $g^{2}(\tau)$ is given by PhysRevA.41.475
@@
In the above equation, if $g^{2}(\tau)\leq 1$ for all $\tau$, it implies $\langle(\Delta n)^{2}\rangle<\langle N\rangle$, i.e., a sub-Poissonian statistics, indicating non-classicality of light. However, if $g^{2}(\tau)\leq 1$ for some $\tau$, there is no direct relation between the sub-Poissonian photon count statistic and the antibunching of photons. Xou and Mandel PhysRevA.41.475 showed that if a source emits photons in two different nonvacuum modes, both having occupation number $n/2$, but different frequencies $\omega_{1}$ and $\omega_{2}$, the second-order correlation function takes the form